Explain what ideas led Bohr to suggest a new atomic model; and state Bohr postulates. Explain also, in words, what the correspondence principle is in atomic physics
1
Expert's answer
2011-06-22T13:50:03-0400
Niels Bohr proposed, in 1913, what is now called the Bohr model of the atom. He suggested that electrons could only have certain classical motions:
1. The electrons can only travel in certain orbits: at a certain discrete set of distances from the nucleus with specific energies.
2. The electrons of an atom revolve around the nucleus in orbits. These orbits are associated with definite energies and are also called energy shells or energy levels. Thus, the electrons do not continuously lose energy as they travel in a particular orbit. They can only gain and lose energy by jumping from one allowed orbit to another, absorbing or emitting electromagnetic radiation with a frequency ν determined by the energy difference of the levels according to the Planck relation:
∆E=E_2-E_1=hν, where h is Planck's constant.
3. The angular momentum L is restricted to be an integer multiple of a fixed unit:
L=nh/2π=nħ
where n = 1, 2, 3, ... is called the principal quantum number, and ħ = h/2π. The lowest value of n is 1; this gives a smallest possible orbital radius of 0.0529 nm known as the Bohr radius.

Bohr's condition, that the angular momentum is an integer multiple of ħ was later reinterpreted by de Broglie as a standing wave condition: the electron is described by a wave and a whole number of wavelengths must fit along the circumference of the electron's orbit:

nλ=2πr.

Principle of correspondence in atomic physics: all further models of atom must include previous+some new predictions.